OpenStudy (anonymous):

A system of equations is shown below: y = 8x − 2 y = 9x − 7 What is the solution to the system of equations? (−5, 38) (−5, −38) (5, 38) (5, −38)

OpenStudy (anonymous):

set the two equations equal to each other

OpenStudy (anonymous):

I don't know how to to do that

OpenStudy (anonymous):

y = 8x − 2 ; y = 9x − 7 8x-2=9x-7 x=5 y=8(5)-2 y=40-2 y=38

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

@Deeezzzz

OpenStudy (0_0youscareme):

C.

OpenStudy (0_0youscareme):

A has a negative sign so itz incorrect